OpenStudy (henrietepurina):

A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2 and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (henrietepurina):

Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (anonymous):

@BOSSKID r u gonna get this?

OpenStudy (anonymous):

If you can't I will.

OpenStudy (anonymous):

The slope-intercept form is y=mx+b You are given the equation 2x+3y=1200,solve for y y=mx+b 2x+3y=1200 3y=-2x+1200 (3/3)y= (-2/3)x+(1200/3) y=(-2/3)x+400 y=(-2/3)x+400 is in this formula y=mx+b (the slope of y=mx+b is the m value, and the y-intercept is the b value) So what is the equation?

OpenStudy (anonymous):

Actualy it's supposed to be written in this form --> y=-2/3X 400 So yes you're correct :)

OpenStudy (anonymous):

*y=(-2/3)x+400

OpenStudy (henrietepurina):

thanks

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

Do you need help?

OpenStudy (henrietepurina):

I need a graph or a scanned graph

OpenStudy (anonymous):

@luvalpaca

OpenStudy (anonymous):

im here

OpenStudy (henrietepurina):

I need one not an equation

OpenStudy (anonymous):

@Cpt.Sparz

OpenStudy (anonymous):

Yes?

OpenStudy (henrietepurina):

I need a graph for my question

OpenStudy (anonymous):

look up scanned graph upward in this question

OpenStudy (anonymous):

You can look it up on google. Or watch a Khan Academy video.

OpenStudy (anonymous):

Type equation into here and it will make a make-shift graph for you http://www.wolframalpha.com/

OpenStudy (anonymous):

^

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

Use @BOSSKID equation

OpenStudy (henrietepurina):

Thankss Guys

OpenStudy (anonymous):

yw :)

OpenStudy (henrietepurina):

The wolfram alpha thing didnt work plzz help